OpenStudy (anonymous):

f a polynomial function f(x) has roots 1+ square root of 2 and -3, what must be a factor of f(x)?

OpenStudy (mathstudent55):

Unless there is more information given, such as the polynomial must have rational coefficients, this question cannot be answered.

OpenStudy (mathstudent55):

No. sorry. I take that back. i misread it.

OpenStudy (mathstudent55):

A polynomial function with roots, a, b, c, ... has factors (x - a), (x - b), (x - c), ...

OpenStudy (anonymous):

There are answer choices, do you want me to send them too?

OpenStudy (mathstudent55):

In other words, x minus each root is a factor.

OpenStudy (mathstudent55):

Since the roots are \(1 + \sqrt 2\) and -3, x minus each of those is a factor.

OpenStudy (mathstudent55):

Yes, what are the choices?

OpenStudy (anonymous):

One moment.

OpenStudy (mathstudent55):

ok

OpenStudy (anonymous):

Did it send?

OpenStudy (mathstudent55):

Ok. Notice the roots you are given. Let's look at \(1 + \sqrt 2\). One factor is x minus that square root. That would be \(x - (1 + \sqrt 2) \) This does not look like any of the choices, but we can work a little on it. Let's distribute the negative outside the parentheses. \(x - (1 + \sqrt 2)\) \(= x - 1 - \sqrt 2\) This is still not like any choice. Does it mention anywhere in the section where this problem comes from that the polynomials must have rational coefficients?

OpenStudy (anonymous):

No, it just polynomial with real roots. It does not give any other details in the question.